# State Frequency Estimation for Anomaly Detection

### v1

@@ V-B SEQUENT’s Anomaly Detection Pipeline
Computing Root Causes. An advantage of using a state machine to model sequential behavior is the ability to follow and analyze the path that an arbitrary trace $t$ has taken within the model. We leverage this capability to compute a root-cause symbol for $t$. We do so by finding the state with the largest rolling anomaly score visited by $t$ and using the state identifier as the root-cause symbol. It denotes the state that causes the largest anomalous activity in $t$. As an example, let $t=[a,a,b,f,a,b,c]$ be the trace under scrutiny, $q=[1,4,7,8,4,4,4]$ be the sequence of states visited by $t$ and $a=[0.19,0.20,0.18,0.16,0.50,0.60,0.65]$ be the rolling anomaly scores computed for $q$. State 4 has the largest score (0.65) and thus ”4” is the root-cause symbol for $t$.
@@
VI Experiment Setup
VI-A Dataset Selection
@@
The first dataset consists of NetFlows collected from a Kubernetes cluster[9] and used in the study conducted by Cao et al. [10]. It contains benign flows generated by real-life users and malicious flows generated by three attack scenarios. The three scenarios are constructed using a specialized threat matrix of the MITRE ATT&CK framework, designed explicitly for Kubernetes [10]. Each attack is fragmented into various steps that an adversary has to execute (e.g., initial access, execution, discovery, etc.). The three attack scenarios cover attacks such as (1) exploiting a vulnerability to discover secrets and using them to do privilege escalation, (2) executing a denial of service (DoS) attack, and (3) exploiting a vulnerability that allows the deployment of malware.
The second dataset is a well-known NetFlow dataset consisting of a large collection of realistic botnet network traffic mixed with normal and background traffic. Specifically, there are thirteen captures of various botnet samples. Each comprises various actions and protocols exhibited by the deployed malware [11]. The captures cover a wide range of attacks, e.g., distributed DoS, port scans, Click-Fraud, and sending spam. The benign and background flows were collected from normal machines deployed on a university network, and malicious flows were collected from virtual machines that were running the different botnets. Each virtual machine was bridged onto the university network to collect flows for the final dataset.
The third and final dataset consists of 100 GB of NetFlows collected from a Spanish ISP. The benign flow data was collected from real-life clients of the ISP, and the malicious flow data was collected from five machines that were used to launch the attacks. The attacks were not launched against any clients but rather five machines that were set up as victim machines [17]. The dataset mostly covers various types of DoS and scanning attacks. Furthermore, it also covers attacks in which spam emails are sent and botnet traffic. The botnet traffic was collected from the CTU-13 dataset and was modified before they were inserted into this dataset. The modifications consist of matching the IP addresses and timestamps of the flows to the ones used in their setup.
VI-B Creating Train/Test Splits
@@
The AssureMOSS dataset was collected over a two-hour period, during which no attack scenarios were initiated within the first hour. We leverage the first hour as our train split and the second hour as our test split. This provides us with 120 thousand flows for training and 275 thousand flows for testing.
As for the CTU-13 dataset, we follow the split presented by its authors [11]: scenarios 3,4,5,7,10,11,12 and 13 for training and scenarios 1,2,6,8,9 for test. The training scenarios contain malicious flows, so we only use benign flows for training. Moreover, using all data from each scenario will lead to an overly large training and test set. To prevent this, we sample a subset of flows from each scenario, resulting in 640 thousand flows for training and 522 thousand flows for test.
For the UGR-16 dataset, we follow the split presented by its authors [17]: data from March to June are used as training data, and data from July and August are used as test data. As this dataset contains, by far, the largest amount of NetFlow data, we use the same sampling technique as with CTU-13 to create our train and test set. This yields a total of 600 thousand flows as training data and 664 thousand flows as test data.
VI-C Methods for Comparisons
@@
From Literature. The AssureMOSS dataset was recently published by Cao et al. and remains relatively unexplored in the context of evaluating other unsupervised anomaly detection methods. Cao et al. conducted a preliminary evaluation comparing their approach against Isolation Forest and Local Outlier Factor [10].
@@
Regarding the CTU-13 dataset, Zoppi et al. evaluated various unsupervised methods for identifying network anomalies [31, 32]. Specifically, the authors evaluated the efficacy of Isolation Forest and Fast Angle-Based Outlier Detection (FastABOD). Additionally, Grov et al. evaluated the efficacy of Markov Chains in detecting botnets [12]. Pellegrino et al. also conducted a study in which they evaluated the efficacy of (probabilistic) state machines and Bigram models (which are essentially Markov Chains) in detecting botnets [24].
@@
For the UGR-16 dataset, Zoppi et al. evaluated the use of two different methods to detect network attacks [31, 32]. However, due to their inability to handle the size of our train/test splits, we dropped them for comparison. Nguyen et al. utilized this dataset to evaluate the efficacy of GEE. Similar to Kitsune [19], they use a state-of-the-art deep learning technique to detect anomalies. However, GEE explicitly trains a single VAE instead of an ensemble of AE.
@@
We select the following from the compiled list of methods to compare against SEQUENT: Isolation Forest, FastABOD, Markov Chain, (probabilistic) state machines, and GEE. As for the implementation of the selected methods, we use Isolation Forest from the scikit-learn library [8], FastABOD from the PyOD library [30], Flexfringe to learn a Markov Chain and (probabilistic) state machine, and the implementation of GEE written by Wong Mun Hou [29]. For each method, except for the Markov Chain, we use the same parameters reported by their authors to train a model on the corresponding dataset. In cases where no parameter information is provided, we make our best effort to train a model by following the reported training steps. We chose to learn a Markov Chain using FlexFringe as we employ a distinct set of features, and learning a Markov Chain follows a standard procedure.
@@
Additional Methods. We included additional methods beyond the aforementioned ones to compare against SEQUENT. One of these methods is the straightforward Boxplot Method presented by Arp et al.  [6], as it employs the frequency of data to detect anomalies. For our experiment, we wrote our implementation of the Boxplot Method following the description provided by the authors. Zoppi et al. also used Histogram-Based Outlier Detection (HBOD) to detect network anomalies. Although not applied to any of the selected datasets, it utilizes the frequency of bins to detect anomalies. This serves as a nice alternative baseline in our comparison. We use PyOD’s implementation of HBOD. Finally, we included Kitsune as it is a state-of-the-art unsupervised ADS. We use the implementation provided by the authors [20]. As it was designed to run at the packet level, we have to make minimal changes for it to work at the flow level. We recognize that this might negatively impact its performance, but we believe that it is beneficial to have this as a baseline for assessing SEQUENT’s usefulness.
VI-D Feature Selection & FlexFringe Setup
@@
As selected datasets do not report the same flow features, we are required to select a subset of features shared among the selected dataset. The features shared between the selected datasets are timestamp, duration, protocol, number of bytes sent, number of packets sent, source and destination IP addresses, and source and destination ports. From this subset of features, we select the duration, protocol, number of bytes sent, and the number of packets sent to generate traces for SEQUENT. Port numbers are not considered within SEQUENT as they are spurious features within the CTU-13 dataset [22]. We discretize all the selected features except for protocol, as it is already a categorical feature.
@@
We configure FlexFringe to sort the flows based on the source and destination IP pairs as we want to learn a state machine from connection data. Furthermore, we configure FlexFringe to use a sliding window of length 10 for the creation of traces. By default, FlexFringe stops replaying a trace if it reaches a state and no outgoing transition exists for the next event in the trace. We, therefore, configure FlexFringe to allow the replay to reset back to the root state and continue with the remaining events. This enables us to compute the visit frequencies of the states for the complete trace. Our FlexFringe configuration settings are available on our repository [4].
VII Evaluation Results
(a) AssureMOSS dataset.
(b) CTU-13 dataset.
(c) UGR-16 dataset.
Figure 4: ROC Curves of each model across the three selected datasets.
VII-A Effectiveness of SEQUENT
@@
We report SEQUENT’s efficacy in detecting various network anomalies on the selected datasets. We use the Area Under Curve (AUC) score as the comparison metric between the various types of models. We deem this a suitable metric for comparison as each method generates anomaly scores for different traces, and we are uncertain about the optimal classification threshold that would yield the most favorable outcomes. Besides the AUC scores, we also report the Receiver Operator Characteristics (ROC) curves computed for each model. The AUC scores and ROC curves on each dataset are reported in Table I and Figure 4, respectively. These are average performance results computed from 10 independent runs. All models were trained on the same features considered by SEQUENT. We also include models trained on the original feature set used in the original study. These are marked with the $\dagger$-symbol in the table. Furthermore, FF-MC denotes the Markov Chain learned using the FlexFringe framework and FF-SM the (probabilistic) state machine. The reported results demonstrate that SEQUENT consistently outperforms all other models in detecting network anomalies on each respective dataset. Moreover, the results also suggest that SEQUENT is generalizable over different network architectures.
Model AssureMOSS CTU-13 UGR-16 Isolation Forest 0.140 0.473 0.426 Isolation Forest - 0.672 - FastABOD 0.229 0.492 0.576 FastABOD - 0.688 - HBOD 0.195 0.357 0.326 HBOD - 0.589 - Boxplot Method 0.436 0.410 0.738 FF-MC 0.067 0.232 0.392 GEE 0.335 0.473 0.505 GEE - - 0.683 Kitsune 0.371 0.5 0.621 FF-SM 0.041 0.142 0.257 SEQUENT 0.978 0.874 0.933
@@
TABLE I: Average AUC of each model across all datasets.
@@
Figure 5: Simplified state machine model learned from CTU-13 dataset.
@@
Figure 6: Zoomed-in state machine of Figure 5, showing the top three anomalous states (47, 62 and 31) presented in Table II.
VII-B Example Analysis Using Root Causes
@@
One of SEQUENT’s advantages is the ability to group anomalies using the generated root-cause symbols. This enables an analyst to reduce the time spent analyzing false positives by filtering out symbols deemed to be false positives. We provide an illustrative example to demonstrate how these symbols can aid with the analysis of anomalies. Let’s say that an analyst has set an arbitrary threshold on the (rolling) anomaly scores to select ”interesting” alerts for further analysis. The analyst then groups the selected traces based on their root-cause symbols and computes the size of each group. Table II presents the top 3 largest groups of anomalies computed from the CTU-13 dataset. Additionally, Figure 5 illustrates the simplified state machine learned by SEQUENT. The analyst decides to first analyze group ”62” by investigating the corresponding (starting) flows collected and linked by SEQUENT (see Table III). The most anomalous flows are shown at the top. SEQUENT deems these flows to be anomalous due to the significant increase in network behavior that coincidentally correlates with the use of port 25, which is used for Simple Mail Transfer Protocol (SMTP). The analyst deems these flows to be suspicious due to the destination port used in the communication and marks all traces with the root-cause symbol ”62” as malicious. Examining the actual labels of these flows, they correspond to the malicious traffic produced by the Neris botnet. This malware specifically sends spam mail [11], which explains the use of the SMTP port (25).
Root-Cause Symbol Size 62 18646 47 3182 31 2556
@@
TABLE II: Anomaly group sizes computed using root-cause symbols generated for the CTU-13 dataset
@@
TABLE III: Simplified flows linked to the top 10 anomalous traces for root-cause symbol ”62”.
Following the same approach, the analyst now investigates group ”47” using the linked flows depicted in Table IV. SEQUENT does not know the actual nature behind these flows, but due to the significant increase in this type of network behavior, it deems these flows to be anomalous. Upon inspection, the analyst observes that these are correlated with the use of port 53, which is typically used for domain name resolution (DNS). This significant increase could be related to a DNS flood attack, or it could simply be that this type of network behavior was rarely seen during training. The analyst deems these flows unrelated to an attack and marks all traces grouped under the symbol ”47” as false positives. By investigating only a subset of the linked flows, the analyst can filter out 3182 false positive alerts and reduce the time spent analyzing every alert. Furthermore, the analyst can visualize the part of the model causing the most anomalous activity (see Figure 6) to understand the sequential behavior exhibited by the most anomalous groups of traces.
@@
TABLE IV: Simplified flows linked to the top 10 anomalous traces for root-cause symbol ”47”.
VII-C SEQUENT’s Robustness
An essential characteristic of an effective ADS is its ability to withstand attacks launched by adversaries seeking to evade detection. With various works showing that ML models are susceptible to adversarial attacks [5, 7, 27], it is imperative to evaluate SEQUENT’s efficacy against (potential) adversarial attacks. To so do, we constructed four distinct types of adversarial attacks following the adversary’s capabilities presented in Section III. We strictly consider attacks that an adversary could launch at test time, and we do not consider adversarial attacks such as poisoning of training data. An adversary is only allowed to adapt the features considered by SEQUENT for anomaly detection, and they are allowed to collect benign flows from the targeted network before the start of an attack. The constructed adversarial attacks are defined as follows:

### v2

@@ 17       $StateScore\leftarrow StateScore+log(TotalVisit+1)$
Computing Root Causes. An advantage of using a state machine to model sequential behavior is the ability to follow and analyze the path that an arbitrary trace $t$ takes within the model. We leverage this capability to compute a root-cause symbol for $t$. We do so by finding the state with the largest rolling anomaly score visited by $t$ and using the state identifier as the root-cause symbol. It denotes the state that causes the largest anomalous activity in $t$. As an example, let us use the model shown in Figure 1 as the model that was learned from benign flows. Let $t{=}[a,a,b,a,b,a,a]$ be the trace under scrutiny, $seq{=}[1,1,2,1,2,1,1]$ be the sequence of states visited by $t$ and $a{=}[0.19,0.25,0.18,0.30,0.20,0.35,0.40]$ be the rolling anomaly scores computed for $seq$. State 1 has the highest anomaly score (0.40), making “1” the root-cause symbol for $t$.
@@
6. Experiment Setup
6.1. Dataset Selection
@@
AssureMOSS Dataset. The first dataset consists of NetFlows collected from a Kubernetes cluster(Cao and Blaise, 2022) and used in the study conducted by Cao et al. (Cao et al., 2022). It contains benign flows generated by real-life users and malicious flows generated by three attack scenarios. The three scenarios are constructed using a specialized threat matrix of the MITRE ATT&CK framework, designed explicitly for Kubernetes (Cao et al., 2022). Each attack is fragmented into various steps that an adversary has to execute (e.g., initial access, execution, discovery, etc.). The three attack scenarios cover attacks such as (1) exploiting a vulnerability to discover secrets and using them to do privilege escalation, (2) executing a denial of service (DoS) attack, and (3) exploiting a vulnerability that allows the deployment of malware. Cao et al. created the dataset to investigate the effectiveness of state machines for the aforementioned network attacks. For this reason, we selected this dataset as one of the baselines for our work.
CTU-13 Dataset. The second dataset is a well-known NetFlow dataset consisting of a large collection of realistic botnet network traffic mixed with normal and background traffic. Specifically, there are thirteen captures of various botnet samples. Each comprises various actions and protocols exhibited by the deployed malware (Garcia et al., 2014). The captures cover a wide range of attacks, e.g., distributed DoS, port scans, Click-Fraud, and sending spam. The benign and background flows were collected from normal machines deployed on a university network, and malicious flows were collected from virtual machines that were running the different botnets. Each virtual machine was bridged onto the university network to collect flows for the final dataset. This dataset has been used in multiple prior studies that investigated the effectiveness of state machines in detecting botnet traffic (e.g., Pellegrino et al. (Pellegrino et al., 2017) and Grov et al. (Grov et al., 2019)) and serves as another suitable baseline for our work.
UGR-16 Dataset. The third and final dataset consists of 100 GB of NetFlows collected from a Spanish ISP. The benign flow data was collected from real-life clients of the ISP, and the malicious flow data was collected from five machines that were used to launch the attacks. The attacks were not launched against any clients but rather five machines that were set up as victim machines (Maciá-Fernández et al., 2018). The dataset mostly covers various types of DoS and scanning attacks. Furthermore, it also covers attacks in which spam emails are sent and botnet traffic. The botnet traffic was collected from the CTU-13 dataset and was modified before they were inserted into this dataset. The modifications consist of matching the IP addresses and timestamps of the flows to the ones used in their setup. We selected this dataset as one of our baseline datasets because it contains a wide range of network behaviors produced by various users.
6.2. Creating Train/Test Splits
@@
The AssureMOSS dataset was collected over a two-hour period, during which no attack scenarios were initiated in the first hour. In the study conducted by Cao et al., the first hour was used as the training data because it contains only benign flows, while the second hour was used as the test data. Because SEQUENT also learns a state machine from benign flows for anomaly detection, we adopt the same train/test split as used by Cao et al. This results in a training set containing 120,000 flows and a test set containing 275,000 flows.
As for the CTU-13 dataset, we follow the split presented by its authors (Garcia et al., 2014): scenarios 3,4,5,7,10,11,12 and 13 for training and scenarios 1,2,6,8,9 for test. The training scenarios contain benign and malicious flows, so we extracted only benign flows for training. Additionally, using all data from each scenario will lead to an overly large training and test set. To address this, we sampled a subset of flows from each scenario, resulting in 640,000 flows for training and 522,000 flows for test.
For the UGR-16 dataset, we follow the split presented by its authors (Maciá-Fernández et al., 2018): data from March to June are used as training data, while data from July and August are used as test data. As this dataset contains, by far, the largest amount of NetFlow data, we use the same sampling technique as with CTU-13 to create our training and test set. This yields a total of 600,000 flows as training data and 664,000 flows as test data.
6.3. Methods for Comparisons
@@
From Literature. The AssureMOSS dataset was recently published by Cao et al. and remains relatively unexplored in the context of evaluating other unsupervised anomaly detection methods. Cao et al. conducted a preliminary evaluation comparing their approach against Isolation Forest and Local Outlier Factor (Cao et al., 2022).
@@
Regarding the CTU-13 dataset, Zoppi et al. evaluated various unsupervised methods for identifying network anomalies (Zoppi et al., 2021a, b). Specifically, the authors evaluated the efficacy of Isolation Forest and Fast Angle-Based Outlier Detection (FastABOD). Additionally, Grov et al. evaluated the efficacy of Markov Chains in detecting botnets (Grov et al., 2019). Pellegrino et al. also conducted a study in which they evaluated the efficacy of (probabilistic) state machines and Bigram models (which are essentially Markov Chains) in detecting botnets (Pellegrino et al., 2017).
@@
For the UGR-16 dataset, Zoppi et al. evaluated the use of two different methods to detect network attacks (Zoppi et al., 2021a, b). However, due to their inability to handle the size of our train/test splits, we dropped them for comparison. Nguyen et al. utilized this dataset to evaluate the efficacy of GEE. Similar to Kitsune (Mirsky et al., 2018), they use a state-of-the-art deep learning technique to detect anomalies. However, GEE explicitly trains a single VAE instead of an ensemble of AE.
@@
We select the following from the compiled list of methods to compare against SEQUENT: Isolation Forest, FastABOD, Markov Chain, (probabilistic) state machines, and GEE. As for the implementation of the selected methods, we use Isolation Forest from the scikit-learn library (Buitinck et al., 2024), FastABOD from the PyOD library (Zhao, 2024), Flexfringe to learn a Markov Chain and (probabilistic) state machine, and the implementation of GEE written by Wong Mun Hou (Wong, 2024). For each method, except for the Markov Chain, we use the same parameters reported by their authors to train a model on the corresponding dataset. In cases where no parameter information is provided, we make our best effort to train a model by following the reported training steps. We chose to learn a Markov Chain using FlexFringe as we employ a distinct set of features, and learning a Markov Chain follows a standard procedure.
@@
Additional Methods. We included additional methods beyond those previously mentioned to compare against SEQUENT. One of these methods is the straightforward Boxplot Method presented by Arp et al.  (Arp et al., 2022), which employs the frequency of data to detect anomalies. Furthermore, it enables us to determine whether a simple method is already sufficient to detect anomalies in our constructed test sets or if a more complex approach is required. For our experiment, we wrote our implementation of the Boxplot Method following the description provided by the authors. Zoppi et al. also used Histogram-Based Outlier Detection (HBOD) to detect network anomalies. Although not applied to any of the selected datasets, it utilizes the frequency of bins to detect anomalies. This serves as a suitable alternative baseline in our comparison. We use PyOD’s implementation of HBOD. Finally, we included Kitsune as it is a state-of-the-art unsupervised ADS. We use the implementation provided by the authors (Mirsky et al., 2024). As it was designed to run at the packet level, we have to make minimal changes for it to work at the flow level. We recognize that this might negatively impact its performance, but we believe that it is beneficial to have this as a baseline for assessing SEQUENT’s usefulness.
6.4. Feature Selection & FlexFringe Setup
@@
As selected datasets do not report the same flow features, we are required to select a subset of features shared among the selected dataset. The features shared between the selected datasets are timestamp, duration, protocol, number of bytes sent, number of packets sent, source and destination IP addresses, and source and destination ports. From this subset of features, we select the duration, protocol, number of bytes sent, and the number of packets sent to generate traces for SEQUENT. Port numbers are not considered within SEQUENT as they are spurious features within the CTU-13 dataset (Nadeem et al., 2023). We discretize all the selected features except for protocol, as it is already a categorical feature.
@@
We configure FlexFringe to sort the flows based on the source and destination IP pairs as we want to learn a state machine from connection data. Furthermore, we configure FlexFringe to use a sliding window of length ten to generate traces for the training and test sets. By default, FlexFringe stops replaying a trace if it reaches a state and no outgoing transition exists for the next event in the trace. We, therefore, configure FlexFringe to allow the replay to reset back to the root state and continue with the remaining events. This enables us to compute the visit frequencies of the states for the complete trace. Our FlexFringe configuration settings are available on our repository (Cao, 2022b).
7. Evaluation Results
Figure 4. Average ROC curves computed for the AssureMOSS dataset.
Figure 5. Average ROC curves computed for the CTU-13 dataset.
Figure 6. Average ROC curves computed for the UGR-16 dataset.
7.1. Effectiveness of SEQUENT
@@
We report SEQUENT’s efficacy in detecting various network anomalies on the selected datasets. We use the Area Under Curve (AUC) score as the comparison metric between the various types of models. We deem this a suitable metric for comparison as each method generates anomaly scores for different traces, and we are uncertain about the optimal classification threshold that would yield the most favorable outcomes. Besides the AUC scores, we also report the Receiver Operator Characteristics (ROC) curves computed for each model. The AUC scores and ROC curves for each dataset are reported in Table 1, Figure 4, Figure 5, and Figure 6, respectively. These are average performance results computed from 10 independent runs. All models were trained on the same features considered by SEQUENT. We also include models trained on the original feature set used in the original study. These are marked with the $\dagger$-symbol in the table. Furthermore, FF-MC denotes the Markov Chain learned using the FlexFringe framework and FF-SM the (probabilistic) state machine. The reported results demonstrate that SEQUENT consistently outperforms all other models in detecting network anomalies on each respective dataset. Moreover, the results also suggest that SEQUENT is generalizable over different network architectures.
Model AssureMOSS CTU-13 UGR-16
Isolation Forest 0.140 0.473 0.426
Isolation Forest - 0.672 -
FastABOD 0.229 0.492 0.576
FastABOD - 0.688 -
HBOD 0.195 0.357 0.326
HBOD - 0.589 -
Boxplot Method 0.436 0.410 0.738
FF-MC 0.067 0.232 0.392
GEE 0.335 0.473 0.505
GEE - - 0.683
Kitsune 0.371 0.5 0.621
FF-SM 0.041 0.142 0.257
SEQUENT 0.978 0.874 0.933
Table 1. Average AUC of each model across all datasets.
7.2. Runtime Analysis of SEQUENT
To evaluate SEQUENT’s runtime costs, we measured its training times and compared them to those of the other ML models. Additionally, we computed SEQUENT’s average runtime for computing rolling anomaly scores at test time. Table 2 presents the average training times of all ML models across all datasets. Table 3 presents the average time SEQUENT takes to compute a rolling anomaly score for a given state at test time. The runtime averages were obtained from the ten independent runs of our experiments.
From the training times presented in Table 2, we observe that SEQUENT takes approximately half a minute to learn a model from benign flows. While it is not the fastest of all models, it is considerably faster than models such as GEE and FastABOD. In our experiments, we used a GPU to train GEE, whereas SEQUENT does not require a GPU to learn a state machine. Although many other models have considerably lower training times, they are not as effective as SEQUENT in detecting network anomalies. This reflects a trade-off between training time and a model’s detection performance. We believe that spending a bit more time to train SEQUENT can be justified, as it achieves a considerably higher detection performance than the other models. We further support our belief based on SEQUENT’s runtimes in computing rolling anomalies scores across the datasets (see Table 3. SEQUENT computes anomaly scores in microseconds on average, making it a practical solution for network anomaly detection.
Model AssureMOSS CTU-13 UGR-16
Isolation Forest 0.18 ±$\pm~{}0.01$ 1.34 ±$\pm~{}0.11$ 1.00 ±$\pm~{}0.04$
Isolation Forest - 1.26 ±$\pm~{}0.13$
FastABOD 16.13 ±$\pm~{}0.65$ 88.30 ±$\pm~{}1.26$ 82.34 ±$\pm~{}2.83$
FastABOD - 101.29 ±$\pm~{}6.92$
HBOD 0.02 ±$\pm~{}0.0$ 0.25 ±$\pm~{}0.50$ 0.10 ±$\pm~{}0.0$
HBOD - 0.27 ±$\pm~{}0.54$
Boxplot Method 7.24 ±$\pm~{}0.25$ 67.14 ±$\pm~{}0.41$ 100.50 ±$\pm~{}0.53$
FF-MC 1.07 ±$\pm~{}0.03$ 6.51 ±$\pm~{}0.13$ 5.22 ±$\pm~{}0.09$
GEE 99.00 ±$\pm~{}0.10$ 127.44 ±$\pm~{}0.57$ 169.69 ±$\pm~{}0.57$
GEE - - 169.55 ±$\pm~{}0.35$
Kitsune 1.24 ±$\pm~{}0.08$ 6.38 ±$\pm~{}0.26$ 6.06 ±$\pm~{}0.22$
FF-SM 4.02 ±$\pm~{}0.12$ 17.41 ±$\pm~{}1.18$ 22.88 ±$\pm~{}1.53$
SEQUENT 5.11 ±$\pm~{}0.30$ 19.23 ±$\pm~{}0.22$ 27.21 ±$\pm~{}2.55$
@@ 7.2. Runtime Analysis of SEQUENT
Table 2. Average training time (seconds) computed across all datasets.
Dataset Avg. Runtime for Score Computation
AssureMOSS 2.17 ±$\pm~{}3.96$
CTU-13 2.20 ±$\pm~{}5.17$
UGR-16 2.36 ±$\pm~{}3.62$
Table 3. Average runtime (microseconds) for computing rolling anomaly score for a state.
@@
Figure 7. Simplified state machine model learned from CTU-13 dataset.
@@
Figure 8. Zoomed-in state machine of Figure 7, showing the top three anomalous states (47, 62 and 31) presented in Table 4.
7.3. Example Analysis Using Root Causes
@@
One of SEQUENT’s advantages is the ability to group anomalies using the generated root-cause symbols. This enables an analyst to reduce the time spent analyzing false positives by filtering out symbols deemed to be false positives. We provide an illustrative example to demonstrate how these symbols can aid with the analysis of anomalies. Let’s say that an analyst has set an arbitrary threshold on the (rolling) anomaly scores to select ”interesting” alerts for further analysis. The analyst then groups the selected traces based on their root-cause symbols and computes the size of each group. Table 4 presents the top 3 largest groups of anomalies computed from the CTU-13 dataset. Additionally, Figure 7 illustrates the simplified state machine learned by SEQUENT. The analyst decides to first analyze group ”62” by investigating the corresponding (starting) flows collected and linked by SEQUENT (see Table 5). The most anomalous flows are shown at the top. SEQUENT deems these flows to be anomalous due to the significant increase in network behavior that coincidentally correlates with the use of port 25, which is used for Simple Mail Transfer Protocol (SMTP). The analyst deems these flows to be suspicious due to the destination port used in the communication and marks all traces with the root-cause symbol ”62” as malicious. Examining the actual labels of these flows, they correspond to the malicious traffic produced by the Neris botnet. This malware specifically sends spam mail (Garcia et al., 2014), which explains the use of the SMTP port (25).
Root-Cause Symbol Size
62 18646
47 3182
31 2556
1501 2448
72 2244
@@
Table 4. Anomaly group sizes computed using root-cause symbols generated for the CTU-13 dataset
@@
Table 5. Simplified flows linked to the top 10 anomalous traces for root-cause symbol ”62”.
Following the same approach, the analyst now investigates group ”47” using the linked flows depicted in Table 6. SEQUENT does not know the actual nature behind these flows, but due to the significant increase in this type of network behavior, it deems these flows to be anomalous. Upon inspection, the analyst observes that these are correlated with the use of port 53, which is typically used for domain name resolution (DNS). This significant increase could be related to a DNS flood attack, or it could simply be that this type of network behavior was rarely seen during training. The analyst deems these flows unrelated to an attack and marks all traces grouped under the symbol ”47” as false positives. By investigating only a subset of the linked flows, the analyst can filter out 3182 false positive alerts and reduce the time spent analyzing every alert. Furthermore, the analyst can visualize the part of the model causing the most anomalous activity (see Figure 8) to understand the sequential behavior exhibited by the most anomalous groups of traces.
@@
Table 6. Simplified flows linked to the top 10 anomalous traces for root-cause symbol ”47”.
7.4. SEQUENT’s Robustness
An essential characteristic of an effective ADS is its ability to withstand attacks launched by adversaries seeking to evade detection. With various works showing that ML models are susceptible to adversarial attacks (Apruzzese et al., 2019; Biggio et al., 2013; Szegedy et al., 2014), it is imperative to evaluate SEQUENT’s efficacy against (potential) adversarial attacks. To so do, we constructed four distinct types of adversarial attacks following the adversary’s capabilities presented in Section 3. We strictly consider attacks that an adversary could launch at test time, and we do not consider adversarial attacks such as poisoning of training data. An adversary is only allowed to adapt the features considered by SEQUENT for anomaly detection, and they are allowed to collect benign flows from the targeted network before the start of an attack. The constructed adversarial attacks are defined as follows: